What is the center and the radius of a circle whose equation is x^2+y^2 -4x +14y +28=0 ?
Answer by lwsshak3(6490)

What is the center and the radius of a circle whose equation is
x^2+y^2 -4x +14y +28=0 ?
complete the square
x^2 -4x+y^2+14y +28=0
(x^2-4x+4)+(y^2+14y+49)=-28+4+49
(x-2)^2+(y+7)^2=25
Standard form of equation for a circle:
(x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
For given equation:
center:(2,-7)
r^2=251
r=√25=5